Two hemispheres have volumes in the ratio a^2/b^3. What is the ratio of their surface areas?
Answer by sachi(548)

Two hemispheres have volumes in the ratio a^2/b^3. What is the ratio of their surface areas?
ans:- i think the ratio must be a^3/b^3
then the volume of hemisphere=1/2(4/3pi r^3)
being the coeffiicent 1/2(4/3pi ) is constant the ratio will be simply R^3/r^3
so R^3/r^3= a^3/b^3
or R/r=a/b
so ratio of areas=R^2/r^2=a^2/b^2 ans